OpenStudy (anonymous):

prove the identity: cos(x+y) + cos(x-y) = 2cosxcosy

OpenStudy (anonymous):

cos(x+y) = cosx * cosy - sinx * siny cos(x-y) = cosx * cosy + sinx * siny so cos(x+y) + cos(x-y) = 2 cosx*cosy